# Towards Swift Serverless LLM Cold Starts with ParaServe

### v1

@@ 6.2 Key-Value Cache Migration
After reducing the pipeline parallelism size, we migrate all uncompleted requests to one worker for better performance, while allowing for earlier release of resources occupied by other workers. Since model layers are assigned to different workers, the KV cache of all requests is distributed across workers, necessitating the KV cache migration.
ParaServe performs KV cache migration inside a pipeline parallelism group, where the KV cache of layers is distributed across workers. Before migration, we first stop scheduling of existing requests and wait for all on-the-fly batches return. Next, we query the cache block manager to obtain the blocks that are used by existing requests, and then collect these blocks from all workers using a gather primitive in collective communication. Blocks are gathered to the first worker and placed at different layers, according to which worker it comes from.
The KV cache migration process in ParaServe is highly optimized. First, the whole migration workflow is performed in an individual thread and uses low-priority CUDA streams so that the inference tasks will not be affected. Second, we create multiple CUDA streams when moving data from or to GPUs to utilize the GPUs’ high parallelism. Furthermore, the data transmission is performed in a streaming manner. On the target worker, once a chunk of tensors arrives, it is instantly loaded to GPU in a separate CUDA stream. On other workers, once a chunk of tensors is loaded from GPU to host, we immediately send them to the network.
@@ 7 Implementation
We implement ParaServe on top of vLLM [44]. We modify vLLM and add around 3200 lines of code in C++ and Python. We also develop a serverless LLM serving framework with around 3000 lines of code in Python. The serving framework performs request routing and automatically scales number of vLLM workers for each model to match current loads. To improve the cold start performance, we run inference in eager mode that does not use CUDA graphs.
Parameter Manager. We develop parameter manager as a TorchScript [54] C++ library. The entry Python program in container first loads the library and triggers the loading thread, which accesses the shared memory created by model prefetcher and loads the model into GPU. The vLLM engine later queries the loading thread and obtains loaded tensors in order.
Engine Initialization Optimizations. We perform several optimizations to the initialization process of vLLM. First, vLLM allocates CPU key-value cache as swapping space during initialization, costing non-negligible time. Thus, we postpone the CPU cache initialization phase. Second, vLLM runs a random input for the model upon loading completes to profile the amount of free memory during inference and calculate the number of key-value cache blocks. We calculate the amount of free memory based on number of model layers and shapes of intermediate results, thereby skipping the profiling phase. Third, vLLM first initializes the dummy model on CPU, and then substitute existing tensors with newly loaded tensors, and finally transmit these tensors to GPU. We allow the engine to directly employ GPU tensors provided by the parameter manager.
@@ 8 Evaluation
In this section, we present experimental results to validate the efficiency and effectiveness of ParaServe. Our evaluation demonstrates that ParaServe achieves a substantial reduction in TTFT, outperforming baselines by up to 4.7 (§8.2). In end-to-end experiments, ParaServe improves TTFT SLO attainment by up to 1.74 across various loads and constraints while maintaining minimal TPOT SLO violations (§8.3). Furthermore, our analysis of pipeline consolidation reveals that scaling down reduces end-to-end generation time by 2.29, while scaling up lowers the average TTFT under bursty requests by 1.87 (§8.4). Lastly, we wrap up with brownfield results and show that ParaServe achieves an average TTFT reduction of 2.6 in production environment (§8.5).
@@ 8.1 Experiments Setup
Testbed. We have two testbeds: (i) a GPU cluster that contains 4 A10 servers and 4 V100 servers. Each A10 server has a single NVIDIA A10 GPU and 188GB memory, while each V100 server has four NVIDIA V100 GPUs and 368GB memory. The network bandwidth per server is 16Gbps. (ii) a GPU cluster consisting of 2 servers, each equipped with four NVIDIA A10 GPUs, and another 4 servers, each equipped with four NVIDIA V100 GPUs. The A10 servers have 752GB memory and a network bandwidth of 64Gbps per server, while the V100 servers have 368GB memory and a network bandwidth of 16Gbps per server. Both testbeds are connected to a remote storage that hosts the models. The remote storage has sufficient network capacity and the model fetching bandwidth is constrained only by the bandwidth of GPU servers. Since our GPU servers do not have NVLink [55], all evaluated models are able to reside in single card to avoid performance degradation.
Baselines. We compare ParaServe against two baselines:
Serverless vLLM. vLLM [44] is a LLM serving engine and we equip it with the same serverless LLM serving framework with ParaServe. During a cold start, the scheduler iterates through all GPU servers and selects the first one with sufficient GPU resources to host the worker. For a fair comparison, we disable CUDA graph in serverless vLLM.
ServerlessLLM [27]. ServerlessLLM is a state-of-the-art serverless LLM serving system that reduces the cold start latency by loading-optimized checkpoint and caching. Due to the lack of high-speed SSDs in our testbed, we can only allow ServerlessLLM to leverage all free memory on servers to cache models. We use Kubernetes [56] to deploy ServerlessLLM and the containers are created at the beginning, so there will be no container creation overhead during serving. We also configure ServerlessLLM to use vLLM [44] as the serving backend and the CUDA graph is disabled.
@@ 8.2 Cold Start Latency
We first evaluate the cold start latency of different systems on testbed (i). We conduct experiments on both V100 and A10 GPUs and configure ParaServe to use a parallelism size of 4. For ServerlessLLM, we measure its performance with and without model caching. Figure 8 illustrates the time to first token for different models. The results demonstrate that ParaServe achieves the shortest cold start latency for all models. Specifically, ParaServe reduces cold start latency by up to 4.7 compared to serverless vLLM and by up to 3.1 to ServerlessLLM. This performance improvement is attributed to parallelized model fetching and worker-level optimizations.
Furthermore, when ParaServe is constrained to use single worker, it also obtains better performance than ServerlessLLM. For smaller models, such as OPT-6.7B, ParaServe with single worker even achieves a short cold start latency than ServerlessLLM with cached models. This improvement primarily comes from ParaServe’s worker-level overlapping strategies and the optimizations applied to vLLM’s cold-start process.
Performance Breakdown. To better comprehend the performance of ParaServe, we conduct a detailed breakdown of techniques employed in ParaServe. Figure 9 illustrates the incremental improvement achieved by applying each proposed technique. Starting with the original vLLM system, we apply the following techniques step-by-step: model prefetching (+Prefetch), streaming loading and implementation optimizations (+Stream), overlapped model loading and library initialization (+Overlap), and parallelized model fetching (+Parallel). The results show that each technique contributes to a reduction in cold start latency and the cumulative effect of all techniques results in a substantial overall improvement.
@@ 8.3 End-to-End Experiments
We further evaluate the effectiveness of ParaServe through comprehensive end-to-end experiments.
Workloads. We choose the Llama2 model series [16] with FP16 precision in the end-to-end experiments. Following prior work [39], we use three typical LLM applications in experiments: chatbot, code completion, and summarization. Requests for these applications are sampled from ShareGPT [57], HumanEval [58] and Longbench [59], respectively. Since there is no available SLO settings for these applications, we derive SLOs based on the performance of warm requests. Specifically, we measure the TTFT and TPOT for warm requests using Llama2-7B and Llama2-13B, with each request containing 1024 input tokens and a batch size of 8. The results are shown in Table 1. The global TTFT SLO is set to five times the TTFT of warm requests, while the TPOT SLO is defined as a stricter constraint, being twice the TPOT of warm requests.
@@
We leverage Microsoft Azure Function Trace [30] to generate workloads, similar to the approach taken by prior works [38, 27]. Models are mapped to functions in the trace using a round-robin approach, and the requests are sampled from the trace using a Gamma distribution. We manage the sampling process by controlling the coefficient of variance (CV) and request rate per second (RPS).
@@
Effectiveness under different CVs. Figure 10 demonstrates the TTFT SLO attainment of the systems under different CVs. To assess the effectiveness of cache, we also evaluate ParaServe with caching enabled. The results indicate that as RPS increases, TTFT SLO attainment decreases due to resource lacking during bursty requests. Nevertheless, ParaServe consistently satisfies the majority of TTFT SLO requirements under heavy loads, achieving up to 1.74 higher compared to baselines in all scenarios. This is because ParaServe select appropriate pipeline parallelism size based on user SLOs and distribute cold-start workers to mitigate network contention. In contrast, ServerlessLLM exhibits high SLO violations due to limited effectiveness of caching for long-tail models. Enabling cache in ParaServe further improves the TTFT SLO attainment by up to 1.11, as frequently accessed models benefit from caching.
@@
For TPOT SLO attainment, both ParaServe and baselines achieve over 95% attainment in most scenarios, and more than 90% under all CVs and RPS. This level of performance is sufficient for most use cases.
@@
Effectiveness under different SLO scales. We evaluate the systems under different TTFT and TPOT SLOs by adjusting a global SLO scaling parameter, with CV fixed at 8. As shown in Figure 11(a), under tight SLO conditions, all systems experience significant SLO violations, as the time granted for preparing a cold-start worker is exceedingly limited. In such cases, the TTFT SLO attainment of all systems is capped at 63%. However, ParaServe still outperforms baselines, as its faster worker initialization reduces the waiting time for subsequent requests, even if the first request violates SLO. Figure 11(b) demonstrates the performance under looser SLO conditions. ParaServe achieves up to 1.52 improvement in TTFT SLO attainment compared to baselines, with caching further enhancing this improvement to 1.58.
@@
Application Analysis. We also analyze the performance of systems for different applications. Figure 13 illustrates the TTFT SLO attainment of chatbot, code completion, and summarization applications under CV=8 and RPS=0.6. First, the results show that ParaServe significantly enhances the TTFT SLO attainment for chatbot and code application, with up to 1.61 and 1.70 improvement, respectively. The code application has lower SLO attainment compared to others because code completion tasks (i.e., requests in HumanEval dataset [58]) have shorter average output length than chat tasks (i.e., requests in ShareGPT dataset [57]) [39]. Therefore, workers for code completion models keep alive for shorter time, leading to more cold starts. Since the summarization application has loose SLO requirements, it meets few SLO violations for all systems.
@@ 8.4 Pipeline Consolidation
To evaluate the effectiveness of pipeline consolidation, we conduct experiments to analyze ParaServe’s performance using two scaling methods. We deploy Llama2-13B on V100 servers in testbed (i) and set the input and maximum output length of a request to 512 tokens.
Scaling down. We demonstrate the benefits of scaling down using a single cold start request that generates 512 tokens. We set pipeline parallelism size to 4 and record the generation time of each token, as shown in Figure 13. With scaling down, the system continues loading the remaining parts of the model in parallel with inference, and migrates the key-value cache of the ongoing request once loading is complete. This allows subsequent tokens to be generated at a faster speed. As a result, scaling down reduces the end-to-end generation time by 2.29, while maintaining almost same inference speeds during the early cold-start period.
Scaling up. We evaluate scaling up by measure ParaServe’s performance under bursty workloads. We set the maximum batch size for each worker to 8 and vary the number of incoming requests. Figure 14(a) illustrates the average TTFT across different loads. The results show that larger pipeline parallelism sizes significantly reduce TTFT, enabling the system to increase throughput earlier. For example, when handling 128 concurrent requests (the maximum load for 16 V100 GPUs), using four workers in a pipeline parallelism group reduces the average TTFT by 1.87. Furthermore, Figure 14(b) indicates that scaling up incurs little inference overhead, with the average TPOT increasing by 1.08-1.19. This increase is attributed to transmission overhead of intermediate results.
@@ 8.5 Brownfield Evaluation
We implement ParaServe in the production environment and evaluate its effectiveness. Specifically, we deploy multiple serverless functions for cold-start instances while incorporating the parameter manager and engine optimizations. Due to security constraints in the production environment, functions cannot establish direct TCP connections with one another. To address this, we leverage a shared object in remote storage to enable inter-worker communication. For the evaluation, we use vLLM to run Llama2-7B on NVIDIA A10 GPUs with 24GB GPU memory. Requests are generated following Microsoft Azure Function Trace [30]. The bandwidth per function is limited to 5Gbps and each instance is granted a grace period of 5 minutes. Figure 15 shows the cold start latency improvement achieved by ParaServe. The results show that ParaServe significantly reduces cold start latency compared to the original vLLM system, with an average cold-start TTFT reduction of 2.6. This substantial performance improvement stems from ParaServe’s parallelized model fetching and optimized worker initialization processes.
9 Discussion
Deploying larger models. ParaServe is applicable to both models that reside in single GPU server as well as models that are deployed in multiple servers with tensor and pipeline parallelism. For models that are distributed across GPU servers, we can improve their cold-start performance through enlarging the pipeline parallelism size. By controlling the degree of parallelism, we reduce cold start latency to satisfy user SLOs. Moreover, worker-level overlapping and pipeline consolidation can also be utilized in this context.
Maximum request capacity of ParaServe. For each cold-start instance, ParaServe allows it to use more workers or GPU resources than needed to reduce cold start latency, probably decreasing the maximum cold-start requests that the system can serve. However, this issue is effectively mitigated in ParaServe. First, increasing pipeline parallelism size actually improves cluster-level resource utilization by enabling finer-grained GPU memory allocation. Second, for users who allocate extra GPU resources to cold-start workers , these resources are quickly released by adopting pipeline consolidation, so the overhead is minimum. Finally, although pipeline parallelism introduces additional host memory usage due to multiplexed Python runtime, this does not block the launching of workers since host memory is usually adequate in GPU servers. Our end-to-end experiments in §8.3 also validate that ParaServe can achieve high SLO attainment under heavy loads, while baselines cannot.

### v2

@@ 6.2 Key-Value Cache Migration
After reducing the pipeline parallelism size, we migrate all uncompleted requests, allowing for earlier release of resources occupied by other workers. Since model layers are assigned to different workers, the KV cache of these requests is distributed across workers, necessitating the KV cache migration.
HydraServe performs KV cache migration inside a pipeline parallelism group, where layers are distributed across workers. Before migration, we first stop scheduling of existing requests and wait for all on-the-fly batches return. Next, we query the cache block manager to obtain the blocks that are used by existing requests, and then collect these blocks from all workers with a gather operation. Blocks are gathered to the worker with whole model and placed at different layers, according to which worker it comes from.
The KV cache migration process in HydraServe is highly optimized. First, the whole migration workflow is performed in an individual thread and uses low-priority CUDA streams so that the inference tasks will not be affected. Second, we create multiple CUDA streams when moving data from or to GPUs to utilize the GPUs’ high parallelism. Furthermore, the data transmission is performed in a streaming manner. On the target worker, once a chunk of tensors arrives, it is instantly loaded to GPU in a separate CUDA stream. On other workers, once a chunk of tensors is loaded from GPU to host, we immediately send them to the network.
@@ 7 Implementation
We implement HydraServe in both testbed environments and our public serverless inference platform. Our implementation is based on vLLM [40], with a modification of around 3200 lines of code in C++ and Python. For testbed deployment, we develop a custom serverless LLM serving framework with around 3000 lines of code in Python. This framework makes scheduling decisions and automatically scales number of vLLM workers for each model to match current loads. The code of HydraServe is open-source and is publicly available at https://github.com/LLMServe/hydraserve.
Instance startup optimizations. We perform several optimizations to the initialization process of vLLM. First, we postpone the allocation of key-value cache swapping space on CPU to reduce cold start latency. Second, vLLM runs a profiling forward to obtain the amount of free memory during inference. Unlike prior work that profiles this information offline [63], we directly calculate the available free memory based on sizes of intermediate results to skip the online profiling phase. Third, as vLLM first initializes the model on CPU, we allow the instance to directly use GPU tensors provided by the parameter manager through overriding tensor metadata.
Support for large models. HydraServe is applicable to both models that reside in a single GPU as well as those deployed in multiple servers with tensor and pipeline parallelism. Similarly, to improve the cold-start performance of models that are distributed across GPUs, we can increase their pipeline parallelism size and implement worker-level overlapping.
@@ 8 Evaluation
In this section, we present experimental results to validate the efficiency and effectiveness of HydraServe. Our evaluation shows that HydraServe achieves a substantial reduction in TTFT, outperforming baselines by 1.7– 4.7 (§8.2). In end-to-end experiments, HydraServe improves TTFT SLO attainment by 1.43–1.74 across various loads and constraints while maintaining minimal TPOT and cost penalties (§8.3). Our analysis of pipeline consolidation reveals that scaling down reduces end-to-end generation time by up to 2.67, while scaling up shortens the average TTFT under bursty requests by 1.87 (§8.4). Lastly, we wrap up with brownfield results and show that HydraServe achieves an average TTFT reduction of 2.6 in production environment (§8.5).
@@ 8.1 Experiments Setup
Testbed. We have two testbeds. (i) A GPU cluster that contains 4 A10 servers and 4 V100 servers. Each A10 server has a single NVIDIA A10 GPU and 188 GB memory, while each V100 server has four NVIDIA V100 GPUs and 368 GB memory. The network bandwidth per server is 16 Gbps. (ii) A GPU cluster consisting of 2 servers, each having four NVIDIA A10 GPUs, and another 4 servers, each having four NVIDIA V100 GPUs. Each A10 server has 752 GB memory and a network bandwidth of 64 Gbps, while each V100 server has 368 GB memory and 16 Gbps bandwidth. Both testbeds are connected to a remote model storage that has sufficient network capacity. Due to the lack of NVLink [65], all evaluated models can reside in a single GPU to avoid performance degradation.
Baselines. We compare HydraServe against two baselines.
Serverless vLLM. vLLM [40] is a LLM serving engine that can serve single model. To serve multiple models, we equip it with the serverless LLM serving framework of HydraServe. During cold starts, the scheduler iterates through all GPU servers and selects the one with sufficient GPU resources to create a new vLLM serving endpoint.
ServerlessLLM. ServerlessLLM [30] is the state-of-the-art serverless LLM serving system that reduces the cold start latency by loading-optimized checkpoint and caching. Due to the lack of high-speed SSDs in our testbeds, we allocate all available server memory for model caching. We deploy ServerlessLLM on Kubernetes [66] and pre-create its containers to eliminate container creation overhead during serving. The system also uses vLLM [40] as the backend.
We exclude systems that accelerate scaling up via worker cooperation [32, 31] from baselines because (1) they are closed-source and (2) peer fetching offers no benefit over remote storage in our environment.
@@ 8.2 Cold Start Latency
We evaluate the cold start latency for systems on testbed (i), while HydraServe is configured at a parallelism size of 4. For ServerlessLLM, we measure its performance with and without model caching. Figure 7 illustrates the TTFT of systems for different models, showing that HydraServe achieves the shortest cold start latency across all models. Specifically, HydraServe reduces cold start latency by 2.1– 4.7 compared to serverless vLLM and 1.7–3.1 compared to ServerlessLLM. This improvement is attributed to parallelized model fetching and our worker-level optimizations.
Even with a single worker, HydraServe also performs better than ServerlessLLM. For smaller models like OPT-6.7B, HydraServe with single worker even achieves shorter cold start latency than ServerlessLLM with cached models. This advantage comes from HydraServe’s worker-level overlapping strategies and optimizations to vLLM’s cold-start process.
Performance breakdown. To better comprehend the performance of HydraServe, we conduct a detailed breakdown of techniques employed in HydraServe. Figure 8 shows the incremental improvement achieved by applying each proposed technique. Starting with the original vLLM system, we apply the following techniques step-by-step: model prefetching (+Prefetch), streaming loading and implementation optimizations (+Stream), overlapped model and library loading (+Overlap), and parallelized model fetching (+Parallel). The results show that each technique contributes to a reduction in cold start latency and the cumulative effect of all techniques results in a substantial overall improvement.
@@ 8.3 End-to-End Experiments
We further evaluate the effectiveness of HydraServe through comprehensive end-to-end experiments.
Workloads. We choose the Llama2 model series [16] with FP16 precision in the end-to-end experiments. Following prior work [55], we use three typical LLM applications in experiments: chatbot, code completion, and summarization. Requests for these applications are from ShareGPT [67], HumanEval [68] and Longbench [69], respectively. Since there is no available SLO settings for these applications, we derive SLOs based on the performance of warm requests. Specifically, we first measure the TTFT and TPOT for warm requests, with each request containing 1024 input tokens and a batch size of 8. The results are shown in Table 2. We then set the global TTFT SLO to five times the TTFT of warm requests, while applying a stricter TPOT SLO at twice the TPOT of warm requests. This TTFT SLO is already quite stringent for serverless LLM serving, considering industrial LLM serving platforms operate with TTFT SLOs as high as 30s [70].
@@
Following prior work [51, 30], we leverage Microsoft Azure Function Trace [35] to generate workloads. Models are mapped to functions in the trace using a round-robin approach, and requests are sampled from the trace with Gamma distribution. We control the sampling process by changing coefficient of variance (CV) and requests per second (RPS).
@@
Effectiveness under different CVs. Figure 9 demonstrates the TTFT SLO attainment of systems under different CVs. To assess the effectiveness of cache, we also evaluate HydraServe with caching enabled. The results indicate that as RPS increases, TTFT SLO attainment decreases due to a lack of resources during bursty requests. Nevertheless, HydraServe consistently satisfies most of TTFT SLO requirements under different loads, achieving 1.43–1.74 higher TTFT SLO attainment compared to baselines in all scenarios. This is because HydraServe selects appropriate pipeline parallelism size based on user SLOs and distributes cold-start workers to mitigate network contention. Although some workers are allocated with additional resources to reduce TPOT, HydraServe promptly reclaims these resources after pipeline consolidation to maintain performance under heavy loads. In contrast, ServerlessLLM exhibits high SLO violations due to limited effectiveness of caching for long-tail models. Enabling caching in HydraServe further improves the TTFT SLO attainment by up to 1.11 due to benefits for hot models.
@@
For TPOT SLO attainment, both HydraServe and baselines achieve over 95% attainment in most scenarios, and more than 90% attainment under all CVs and RPS configurations. This level of performance is sufficient for most use cases. Detailed results are provided in the appendix due to space limitations.
@@
Effectiveness under different SLO scales. We evaluate the systems under different TTFT and TPOT SLOs by adjusting a global SLO scaling parameter, with CV fixed at 8. Figure 10(a) shows that, under tight SLOs, all systems experience significant SLO violations since the time granted for preparing a cold-start worker is exceedingly limited. In such cases, the TTFT SLO attainment of all systems is capped at 63%. However, HydraServe still outperforms baselines, as its faster worker initialization reduces the waiting time for subsequent requests, even if the first request violates SLO. Figure 10(b) demonstrates the performance under looser SLO conditions. HydraServe achieves 1.38–1.52 improvement in TTFT SLO attainment compared to baselines, with caching further enhancing this improvement to 1.49–1.58.
@@
Application analysis. Figure 12 illustrates the TTFT SLO attainment of chatbot, code completion, and summarization applications under CV=8 and RPS=0.6. First, the results show that HydraServe significantly enhances the TTFT SLO attainment for chatbot and code applications, with up to 1.61 and 1.70 improvement, respectively. The code application has lower SLO attainment compared to others because code completion tasks (i.e., requests in HumanEval dataset [68]) have shorter average output length than chat tasks (i.e., requests in ShareGPT dataset [67]) [55]. Therefore, workers for code completion models keep alive for shorter time, leading to more cold starts. Since the summarization application has loose SLOs, it has few SLO violations for all systems.
TPOT and resource usage penalties. As discussed in §4.1, pipeline parallelism increases the worst-case TPOTs, whereas allocating additional resources helps to reduce them. We evaluate HydraServe’s TPOT and resource usage of different models to examine these trade-offs. Figure 13 illustrates the relative TPOT and cost ratios of HydraServe compared to serverless vLLM under CV=8 and RPS=0.6. The cost is proportional to the GPU memory-time product. Regarding TPOT, HydraServe exhibits only a 1.06 average increase compared to serverless vLLM, with most increases occurring in chatbot and code models that have stringent TTFT requirements. The TPOT penalty is modest because HydraServe can quickly merge pipeline groups into single workers, limiting performance degradation to only the first few tokens. For some models, HydraServe achieves shorter TPOT due to GPU performance fluctuations.
Figure 13(b) reveals that, surprisingly, HydraServe consumes lower cost for most models. This is because (1) for models with additional resource consumption, HydraServe can quickly merge pipeline groups, and (2) HydraServe enables faster worker startup, thereby reducing GPU usage during cold starts. On average, HydraServe reduces costs by 1.12 compared to serverless vLLM.
@@ 8.4 Pipeline Consolidation
To evaluate the effectiveness of pipeline consolidation, we measure HydraServe’s performance with two scaling methods. We deploy Llama2-13B on V100 servers in testbed (i) and set the input and output length of each request to 512 tokens.
Scaling down. We demonstrate the benefits of scaling down by presenting the generation time of each token. We use different batch sizes and set pipeline parallelism size to 4. As shown in Figure 12, with scaling down, the system loads the remaining parts of the model in parallel with inference, and migrates the key-value cache of the ongoing request once loading has completed. This allows subsequent tokens to be generated at a faster speed. As a result, scaling down reduces the end-to-end generation time by 1.90–2.67, while maintaining almost same inference speeds during the early cold-start period.
Scaling up. We evaluate scaling up by measuring HydraServe’s performance under bursty workloads. We set the maximum batch size for each worker to 8 and vary the number of incoming requests. Figure 14(a) illustrates the average TTFT across different loads. The results show that larger pipeline parallelism sizes significantly reduce TTFT, enabling the system to increase throughput earlier. For example, when handling 128 concurrent requests (the maximum load for 16 V100 GPUs), using four workers in a pipeline parallelism group reduces the average TTFT by 1.87. Furthermore, Figure 14(b) indicates that scaling up incurs little inference performance overhead, with the average TPOT increasing by 1.08–1.19. This increase is attributed to the transmission overhead of intermediate results.
@@ 8.5 Brownfield Evaluation
We evaluate the prototype of HydraServe in our production environment. In this environment, functions cannot establish direct TCP connections with one another due to security constraints, so we leverage a shared object in remote storage to enable inter-worker communication. In evaluation, we use vLLM to run Llama2-7B on NVIDIA A10 GPUs with 24 GB GPU memory, with requests generated following Microsoft Azure Function Trace [35]. Figure 15 compares the cold start latency of HydraServe to serverless vLLM. The results demonstrate that HydraServe significantly reduces cold start latency versus the original vLLM system, achieving an average 2.6 reduction in cold-start TTFT. This substantial improvement stems from HydraServe’s parallelized model fetching and optimized worker initialization processes.